### v1

@@ 3.4.5 Corpus Callosum
3.4.6 Brain Stem
The brain stem is also geometrically identical between the gyrencephalic and the lissencephalic models and separated from the cortical folds, and experienced higher peak values for all injury metrics compared to other ROIs, except the corpus callosum. The region experienced overall higher strain in the gyrencephalic models than the lissencephalic ones for all three loading directions ($\textit{p}<0.005$).
The bias (mean $\pm$ std. dev.) in peak MPS95 was found to be $0.064\pm 0.051$ ($p<0.005$), in peak MPSR95 to be $11.84\pm 10.73$ $s^{-1}$ ($p<0.005$), and in CSDM15 to be $14\pm 13.78$ ($p<0.05$) (Figure 6f). The normalized bias values are $10.4\pm 8.4\%$, $9.7\pm 9.3\%$ and $14.7\pm 13.9\%$, in MPS95, MPSR95, and CSDM15, respectively. We found no proportional bias in any injury metric.
Figure 6: Bland-Altman analysis on different regions of interest (ROI), including cerebral gray matter, cerebral white matter, cerebellum, deep brain regions, corpus callosum, and brain stem. The mean bias and limits of agreement (mean±1.96×SD) between the gyrencephalic and lissencephalic peak injury metrics: a) MPS95, b) MPSR95, and c) CSDM15. These trends show the gyrencephalic models overpredict the injury metrics compared to the lissencephalic models with high statistical significance in all ROIs, except for the cerebral white matter, which experiences high variability across individuals and loading directions.
4 Discussion
In this study, we aim to determine whether incorporating cortical folds in FE head models is necessary for accurately predicting the risk of mTBI under head rotational acceleration events. We attempt to answer this question by comparing different mTBI injury metrics between the gyrencephalic and lissencephalic FE models of 18 subjects aged 9 - 18 years. The models were used to simulate an idealized concussive event with the rotational acceleration of 10 krad/s2 and rotational velocity of 60 rad/s about each principal axis. In this section, we discuss the effect of cortical folds on the dynamics of the head model that results in injury metrics differences.
4.1 Effect of cortical folds on brain tissue near the skull-brain interface
We find that the gyrencephalic models overall predict higher peak values of all injury metrics under all three loading directions ($p<0.005$). Since a head acceleration event is modeled through providing the skull with a rigid body acceleration, which results in the strain wave originating at the skull and traveling inwards to the brain through the meninges and CSF. High dissipation of the brain tissue causes the lower strains as the wave travels inwards, and therefore higher strains on the cortical surface (Figure 2a-f) [26]. Therefore, the skull-brain interface plays a significant role in the peak strains experienced by the brain tissue and modeling cortical folds directly impacts the skull-brain interface.
The sulci in the gyrencephalic models are filled with CSF, which has significantly lower stiffness (CSF viscosity $\nu=0.001$ Pa s vs brain stiffness $G_{\infty}$ = 6400 kPa). This results in a reduced effective stiffness of the whole structure, contributing to higher strains in the brain tissues in the gyrencephalic models. We also observe an effect of faster strain wave speed in stiffer brain tissue than the significantly more compliant CSF in CSDM15, which rises faster in the lissencephalic model than in the gyrencephalic model (Figure 4).
The CSF also offers a cushioning effect to the brain tissue, with higher thickness of CSF between the cortical gyri and dura resulting in lower strain on the cortical surface (Supplementary Figure S9). The cushioning effect comes from the dissipation during CSF deformation. The CSF layer in the lissencephalic models is uniformly 1 mm thick and the models consistently experience higher strains near the same regions on the outer cortical surface (along the smaller axes of the cross-section ellipse), under all three directions of rotation (Figure 3). On the other hand, the gyrencephalic models experience non-uniform strain concentration regions on the cortical surface, which are also different between subjects. The non-uniform CSF layer, which is also heterogeneous across the subjects [27], is captured in the FE models through modeling of the cortical folds. The low DICE values for PMPS (0.070.430.07-0.43) and for CSDM15 (0.420.700.42-0.70) indicate that the lissencephalic model predicts significantly different injury distribution than the gyrencephalic models and modeling cortical folds becomes necessary when we need to capture the location of the injury.
4.2 Effect of cortical folds on brain tissue near ventricles
We find that the zone of influence of gyri and sulci extends further inside due to interaction with the deep brain cavity, the ventricles. The brain tissue between the ventricles and sulci experiences higher strain and strain rate concentrations, such as near the Sylvian sulcus (Figure 2). This increased strain concentrations at higher depths in gyrencephalic models, where the strain wave reaches later than the cortical surface, contributes to a higher time to peak MPS95 and MPSR95 (Figure 4) than the lissencephalic models. The interaction between sulci and ventricles also influences the strains and strain rates experienced by the regions of the brain that geometrically remain the same between gyrencephalic and lissencephalic brain, such as the corpus callosum and deep brain regions (Section 2.1). The ROIs that experience smallest difference between the gyrencephalic and lissencephalic model includes the cerebellum, which is shielded from the cortical folds by the tentorium. However, the change in whole head dynamics also affects the deformations experienced by the cerebellum and the brain stem.
Our analyses indicate that modeling cortical gyri results in different mTBI injury metrics, both globally as well as locally. The variations in the cortical folds across individuals also result in local strain distribution variations. Therefore, it is important to incorporate cortical folds in the FE model for accurate injury metrics estimate.
4.3 Comparison with other studies
A number of previous studies have investigated the role of cortical folds on mTBI injury metrics under head acceleration events. However, the studies found conflicting results, with some predicting higher injury metrics in gyrencephalic models [11, 12], some in lissencephalic models [9], and some different trends in different metrics [8]. However, since the loading conditions, material properties for brain, CSF, and the meninges, are all different between the studies, a direct comparison was not possible to determine the importance of modeling cortical folds.
In our study, modeling the cortical folds result in higher peak values of MPS95, MPSR95, and CSDM15 in the whole brain with high statistically significance ($p<0.005$), similar to an impact against a rigid wall at 5m/s using a 2D adaptive lagrangian eularian (ALE) model (CSF modeled as eularian fluid) of the brain [12], a 2D experimental study on a gel phantom under drop impact test [11]. While the conclusions are the same in Fagan et al. (2020) [12], a direct comparison is not feasible due to the different loading conditions of linear acceleration compared to rotational acceleration in our study.
However, Ho and Kleiven (2009) found that lissencephalic models experience higher strains under a 5ms, 10 krad/s2 loading [9]. It should be noted that the CSF is modeled as an elastic fluid, and also incorporates the pia mater, modeled as a viscoelastic shell, significantly altering the skull-brain interface. The choice of skull-brain interface modeling is an open question in the field, and we based our choice on recent studies with validated models [5, 23, 22]. However, the choice of skull-brain interface can significantly alter the brain response, and consequently, the contribution of cortical folds on the brain deformation and injury metrics.
Saez et al (2020) found similar strain distributions in 3D models of human (high gyrification), a macaque (lower gyrification), and a mouse (lissencephalic) scaled to the same size, but higher stresses in lissencephalic brain (mouse) as compared to the gyrencephalic (human and macaque) brains under a linear de-acceleration loading [10]. While the CSF material modeling is not reported in the paper, a direct comparison is not possible since the loading is linear acceleration compared to rotational acceleration in our study.
We also find that the regions of high strain differences are adjacent to deep primary sulci closer to the ventricles (Sylvian sulcus, Figure 3). Multiple mesoscale studies find that as the depth of the sulci increases, the base of the sulci experiences higher strain concentrations [15, 14, 13]. However, since the mesh resolution in our FE models is not as fine as in the meso-scale models, we are unable to capture such strain localization.
@@ 4.4 Limitations and Future Works
Since our workflow involved a lot of manual segmentation and corrections of the automated segmentations of the brain and the skull, some variations in the injury metrics could stem from the variations in the segmentations. Future studies should include automated whole workflows that can offer robust segmentation quality control.
@@
Our FE models also lacked mesh adaptivity or high mesh density near sharp corners, such as cortical sulci and ventricles, limiting the ability to capture the strain concentrations at the sulci. Future studies should move towards higher mesh density near the sharp corners.
@@
Appendix A Segmentation Workflow
The T1-structural MRIs were acquired for all the individuals and pre-processed to correct for Gibbs-ringing and bias-field artifact using open source software MRtrix [29] and ANTS [30], respectively. All MRI scans were registered rigidly to the MNI152 atlas using 3D slicer (http://www.slicer.org) [31]. An automated brain segmentation was performed on the enhanced MRIs using the FreeSurfer software package [17]. The skull was added using FSL [32], and additional segments of the sub-arachnoid cerebrospinal fluid (CSF) and 1 mm thick meninges (dura, falx, and tentorium) were added manually using 3D Slicer after combining the FreeSurfer and FSL segmentations. It was ensured that the CSF layer was at least 1 mm thick everywhere. This workflow provided the gyrencephalic (with cortical folds) model segmentations.

### v2

@@ 4.1 Comparison with other studies
Since incorporating the Pia mater in a voxel mesh with a jagged interface was not feasible in our model, we approximately studied the effect of a stiffer skull-brain interface by modeling the sub-arachnoid space with a higher stiffness of 20kPa, which is stiffer than the long-term modulus of the brain (Supplementary Section S3). Another FE head model has used a similarly stiff sub-arachnoid material to account for the trabeculae and other connective tissues in the region [49]. The result of this preliminary analysis shows that as the CSF stiffness increases, the strain experienced by the brain tissue decreases significantly in the gyrencephalic model. The CSF deformation decreases with the CSF stiffness, and therefore, the strain wave transmitted to the brain is lower than in the case of a softer CSF. Since, the sulci in the gyrencephalic models are filled with CSF, the increased stiffness of the skull-brain interface reduces the strain in the gyrencephalic model more significantly as compared to the lissencephalic model, which experiences a smaller decrease in the strain, resulting in higher strains in the lissencephalic model compared to the gyrencephalic one. Unfortunately, there is no consensus on the best approach to simulate the skull–brain interface, and we the choice of skull-brain interface in our model is based on previously validated models [30, 5, 15, 27, 50].
The other important distinction with the Ho and Kleiven study is the injury metric analyzed. The study looked into the 100th percentile strain, which doesn’t remove the effect of numerical artifacts, as compared to the 95th percentile used in our study. It is also worth mentioning that the quality of segmentation that captures the depth of sulci in the model can be affected by the choice of segmentation tools, which can also contribute to different trends.
The finding of our study that the gyrencephalic models experience higher peak injury metrics is corroborated by most mesoscale models-based studies [17, 19, 16]. However, the boundary conditions to capture the impact loading in these studies are not consistent. Most studies apply compression loading to compare the pressure or stress responses [18, 19, 17]. The studies found increasing stress with increasing depth of the sulci. While [17] also analyzed a shear loading case, they only reported stresses. Therefore, a direct comparison with these studies are not feasible. On the other hand, Cloots et al. [16] approximated the impact event using a shear-acceleration profile in the region of interest, derived from the output of a full-head model simulation under linear acceleration. They also found higher strains and stresses in models with sulci than in a homogeneous model and different strains for different sulci geometries, similar to the results of our study. However, since the model only captures sulci, its interaction with the rest of the brain, especially the ventricles can’t be captured in this model. On the other hand, since the mesh resolution in our FE models is not as fine as in the meso-scale models, we are unable to capture the strain localization at the base of the sulci.
Comparing the results of our study with the existing literature, we find that modeling cortical folds affects the spatial distribution and peak values of the strain-based injury metrics under head rotations. However, the magnitude of differences due to modeling cortical folds is affected by the choice of skull-brain interface and brain stiffness; the mode of loading, translational or rotational; and the injury metrics analyzed.
@@ 4.2 Limitations and Future Works
There are several limitations to our study. First, we only analyze a single head acceleration magnitude of 10 krad/s2, which lies towards the higher end of the range of measured impacts. The duration of the loading (10 ms) also synchronizes with the head natural frequency to produce higher strains than other duration [30]. While we anticipate smaller differences in the peak strains between the lissencephalic and gyrencephalic models as the magnitude of peak strains decreases with other loading conditions, given the complex non-linear relationship between loading and peak injury metrics [30], further analysis is required to access the importance of modeling cortical folds in a different range of loading conditions. However, the general results of this study will still apply to the strain distributions affected by the cortical folds.
Our model didn’t distinctly model the Pia mater, the vasculature in the sub-arachnoid space, and the CSF in the head is also not modeled as a fluid. We modeled the space as a very soft solid to account for all the structures. Unfortunately, the best approach to simulate the skull-brain interface remains not unresolved, but will greatly affect the importance of modeling the cortical folds. Future studies should investigate the effect of modeling Pia mater and cerebral vasculature on the strains experienced by the brain.
@@
Given the resolution of our model, if two gyri touch each other, they share common nodes and therefore are not free to detach from each other. Future studies should investigate the effect of sliding between the gyri. Our FE models also lacked mesh adaptivity or high mesh density near sharp corners, such as cortical sulci and ventricles, limiting the ability to capture the strain concentrations at the sulci. These limits the ability of our model to capture tau-protein accumulation patterns seen in CTE [52]. Future studies should move towards higher mesh density near high curvature points.
@@
  • Supplementary Information The supplementary material is available here.